OpenStudy (anonymous):

Give the 3rd and 7th terms using the sequence formula tn = (n – 3)2. Answer 0, 16 0, 8 0, 49 0, 14

OpenStudy (across):

What is so difficult about this one?

OpenStudy (anonymous):

im confused retricepellet

OpenStudy (across):

Why?

OpenStudy (anonymous):

cuz i know math but i dont really know it. I Learn Math For The Moment Not A Life Time

OpenStudy (across):

That's fair enough. You're given \(t_n=(n-3)^2\) and asked to evaluate that thing at \(n=3\) and \(n=7\):\[t_3=(3-3)^2=0^2=0\\t_7=(7-3)^2=4^2=16\]

OpenStudy (anonymous):

thanks